OpenStudy (anonymous):

what is the change in enthalpy for the reaction of 2.00 mol of iron (ii) oxide

OpenStudy (zale101):

By now, i believe you already know what's this formula stands for delta H=kJ/mol

OpenStudy (zale101):

or you can do this formula too mole=mass/molecular mass

OpenStudy (zale101):

Iron(III) oxide or ferric oxide is the inorganic compound with the formula Fe2O3

OpenStudy (anonymous):

Thanks for your reply I also want ot know what is the energy of one mole of photons with a wavelength of 285 nm

OpenStudy (zale101):

Use the energy of a single photon

OpenStudy (zale101):

E = hv v = c / λ

OpenStudy (zale101):

h = 6.626e-34 J s v = 3x10^8 m/s for electromagnetic waves sorry lol, here is what H and V stands for

OpenStudy (zale101):

E = Energy of a wave h =Planck;s constant v = Velocity of the wave

OpenStudy (zale101):

So far so good?

OpenStudy (anonymous):

your must be

OpenStudy (zale101):

not for me lol, i meant for u

OpenStudy (zale101):

are you ok with everything? What part u don't get?

OpenStudy (zale101):

@hajeralam, u there?

OpenStudy (zale101):

One mole of a substance contains 6.02 x 1023 protons

OpenStudy (zale101):

First, do the frequency (f = c/λ) so you can calculate the energy of a single photon (E = hf)

OpenStudy (zale101):

We know the speed of light is (2.998x10^8 m/s) the wavelength (2.85x102)

OpenStudy (zale101):

f = c/λ = (2.998x10^8 m/s) / (2.85x10^2 m) = solve :)

OpenStudy (zale101):

the wavelength (2.85x10^2)

OpenStudy (anonymous):

thanks so much ur a lifesaver

OpenStudy (zale101):

ur welcome ^_^